# A rapid channel for the collisional formation and gravitational wave driven mergers of supermassive black hole seeds at high redshift

### v1

@@ 4 A characteristic GW fingerprint of collisional SMBH seed formation
At $t=10$ Myr, the hierarchical simulations also host a population of BH binaries with $q\sim 0.1$ consisting of one IMBH in the range of ${300}\>\mathrm{M_{\odot}}\lesssim M_{\mathrm{\bullet,1}}\lesssim{1000}\>\mathrm{M_{\odot}}$ and a stellar BH. At the higher IMBH masses up to $\sim{3000}\>\mathrm{M_{\odot}}$, some IMBH-BH binaries have already merged within $10$ Myr (right panel of Fig. 5). Most interestingly, the hierarchical models form binaries of similar mass with $M_{\mathrm{\bullet,1}}\gtrsim{2400}\>\mathrm{M_{\odot}}$ at $t=10$ Myr and a few IMBH-IMBH binaries with $M_{\mathrm{\bullet,1}}\gtrsim{2400}\>\mathrm{M_{\odot}}$ and $q>0.1$ have already merged, two of them with mass ratios near unity. Therefore, in the hierarchical star cluster assembly scenario, multiple IMBHs can end up in the same assembling star cluster and they can also rapidly merge with each other within $10$ Myr of cluster evolution. The binary IMBH mergers are often driven by a third IMBH. This additional feature of the hierarchical assembly scenario leads to more rapid ($<10$ Myr) IMBH mergers than in timescale considerations of two IMBHs per cluster (Amaro-Seoane & Freitag, 2006; Amaro-Seoane et al., 2010; Arca Sedda & Mastrobuono-Battisti, 2019; Rasskazov et al., 2020; Souvaitzis et al., 2025) driven by stellar hardening alone. In addition, triple IMBH interactions in the hierarchical cluster assembly models may also more often result in above near-zero eccentricities at GW mergers.
[…]
@@ 5 Summary
We show that hierarchical massive star cluster formation at low metallicities (${0.01}\>\mathrm{Z_{\odot}}$) can lead to the formation of IMBHs and the rapid GW driven IMBH-IMBH mergers. The N-body models include evolving single stars and initial binary and triple systems as well as post-Newtonian dynamics for BHs. Initial stellar multiplicity leads to earlier stellar mergers, higher collision rates and a factor of 22223333 larger IMBH masses compared to single star models. The hierarchical cluster assembly models frequently result in the formation of close to equal mass IMBH binaries and GW driven IMBH mergers. Each sub-cluster with $M_{\mathrm{cl}}\gtrsim{10^{4}}\>\mathrm{M_{\odot}}$${10^{5}}\>\mathrm{M_{\odot}}$ may provide a single IMBH for the hierarchically assembled cluster, often resulting in more than two IMBHs in the assembled cluster. In general, models with increasing stellar multiplicity lead to more SMBH seeds in the final cluster. In idealised isolated clusters, even initial stellar multiplicity does not result in IMBH-IMBH mergers.
The hierarchical scenario provides a rapid and efficient channel for IMBH formation through stellar collisions and tidal disruption events, their pairing in massive binaries via star cluster mergers, and finally to potential GW mergers driven by stellar hardening and triple IMBH interactions. The entire process can occurs in less than $10$ Myr. If retained in their formation environments, the IMBHs formed through repeated stellar collisions constitute a natural candidate for the seeds of SMBHs. SMBH seeds up to $\sim{6200}\>\mathrm{M_{\odot}}$ ejected by GW recoil kicks often have relatively low velocities, only up to a few times $100$ km/s. While these kick velocities are high enough to unbind the IMBHs from their host clusters but not from their host galaxies. IMBH-IMBH mergers may constitute an important growth channel for the early SMBH seeds without any AGN or TDE signature. Our results suggest that a $M_{\bullet}\gtrsim{1000}\>\mathrm{M_{\odot}}$ GW merger with similar IMBH masses is a characteristic GW fingerprint of star cluster assembly with simultaneous SMBH seed formation at high redshifts. At lower redshifts the IMBH formation and GW mergers may still proceed in dense, low metallicity environments. Besides GWs, tidal disruption events with TDE rates up to a few times $\Gamma\sim 10^{-5}$ yr-1 per assembling star cluster might reveal a population of such low redshift IMBHs in the upcoming extragalactic wide field time domain surveys.
@@ Acknowledgments
The numerical simulations were performed using facilities hosted by the Max Planck Computing and Data Facility (MPCDF), Germany. TN acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence "ORIGINS".

### v2

@@ 4 A characteristic GW fingerprint of collisional SMBH seed formation
At $t=10$ Myr, the hierarchical simulations also host a population of BH binaries with $q\sim 0.1$ consisting of one IMBH in the range of ${300}\>\mathrm{M_{\odot}}\lesssim M_{\mathrm{\bullet,1}}\lesssim{1000}\>\mathrm{M_{\odot}}$ and a stellar BH. At the higher IMBH masses up to $\sim{3000}\>\mathrm{M_{\odot}}$, some IMBH-BH binaries have already merged within $10$ Myr (right panel of Fig. 5). Most interestingly, the hierarchical models form binaries of similar mass with $M_{\mathrm{\bullet,1}}\gtrsim{2400}\>\mathrm{M_{\odot}}$ at $t=10$ Myr and a few IMBH-IMBH binaries with $M_{\mathrm{\bullet,1}}\gtrsim{2400}\>\mathrm{M_{\odot}}$ and $q>0.1$ have already merged, two of them with mass ratios near unity. Therefore, in the hierarchical star cluster assembly scenario, multiple IMBHs can end up in the same assembling star cluster and they can also rapidly merge with each other within $10$ Myr of cluster evolution. The binary IMBH mergers are often driven by a third IMBH. This additional feature of the hierarchical assembly scenario leads to more rapid (<10<10 Myr) IMBH mergers than in timescale considerations of two IMBHs per cluster (Amaro-Seoane & Freitag, 2006; Amaro-Seoane et al., 2010; Arca Sedda & Mastrobuono-Battisti, 2019; Rasskazov et al., 2020; Souvaitzis et al., 2025) driven by stellar hardening alone. In addition, triple IMBH interactions in the hierarchical cluster assembly models may also more often result in above near-zero eccentricities at GW mergers.
Third-generation ground-based GW detectors like the Einstein Telescope and the Cosmic Explorer will be sensitive to IMBH-IMBH mergers with masses $M_{\mathrm{bin}}=M_{\mathrm{\bullet,1}}+M_{\mathrm{\bullet,2}}\lesssim{10^{3}}\>\mathrm{M_{\odot}}$ at $z<10$ (Reali et al., 2024; Abac et al., 2025). The $M_{\mathrm{bin}}\lesssim{10^{3}}\>\mathrm{M_{\odot}}$ $z<10$ GW mergers most probably occur too late to belong the parent population of SMBHs powering the z=6z=677 active galactic nuclei (Fan et al., 2023; Inayoshi et al., 2020), but can nevertheless constrain the mass function of the seeds that fail to grow at high redshifts. The sensitivity of LISA for equal-mass BH binaries at high redshifts rapidly increases from $M_{\mathrm{bin}}={10^{3}}\>\mathrm{M_{\odot}}$ towards $M_{\mathrm{bin}}={10^{4}}\>\mathrm{M_{\odot}}$ (Amaro-Seoane et al., 2012). Still, only our most massive IMBH binaries could be detectable with LISA at $z>10$. Overall, we expect that IMBH binaries with $M_{\mathrm{bin}}\gtrsim{10^{4}}\>\mathrm{M_{\odot}}$ formed during hierarchical assembly of clusters with $M_{\mathrm{cl}}\gtrsim$ a few times ${10^{6}}\>\mathrm{M_{\odot}}$ will be promising GW targets for LISA high redshifts. Lower-mass ($\lesssim{10^{3}}\>\mathrm{M_{\odot}}$) $z>10$ IMBH mergers will be observed using GW space observatories with shorter interferometer arms such as TianQin (Luo et al., 2016).
@@ 5 Summary
We show that hierarchical massive star cluster formation at low metallicities (${0.01}\>\mathrm{Z_{\odot}}$) can lead to the formation of IMBHs and the rapid GW driven IMBH-IMBH mergers. The N-body models include evolving single stars and initial binary and triple systems as well as post-Newtonian dynamics for BHs. Initial stellar multiplicity leads to earlier stellar mergers, higher collision rates and a factor of 2233 larger IMBH masses compared to single star models. The hierarchical cluster assembly models frequently result in the formation of close to equal mass IMBH binaries and GW driven IMBH mergers. Each sub-cluster with $M_{\mathrm{cl}}\gtrsim{10^{4}}\>\mathrm{M_{\odot}}$${10^{5}}\>\mathrm{M_{\odot}}$ may provide a single IMBH for the hierarchically assembled cluster, often resulting in more than two IMBHs in the assembled cluster. In general, models with increasing stellar multiplicity lead to more SMBH seeds in the final cluster. In idealised isolated clusters, even initial stellar multiplicity does not result in IMBH-IMBH mergers.
The hierarchical scenario provides a rapid and efficient channel for IMBH formation through stellar collisions and tidal disruption events, their pairing in massive binaries via star cluster mergers, and finally to potential GW mergers driven by stellar hardening and triple IMBH interactions. The entire process can occurs in less than $10$ Myr. If retained in their formation environments, the IMBHs formed through repeated stellar collisions constitute a natural candidate for the seeds of SMBHs. SMBH seeds up to $\sim{6200}\>\mathrm{M_{\odot}}$ ejected by GW recoil kicks often have relatively low velocities, only up to a few times $100$ km/s. While these kick velocities are high enough to unbind the IMBHs from their host clusters but not from their host galaxies. IMBH-IMBH mergers may constitute an important growth channel for the early SMBH seeds without any AGN or TDE signature. Our results suggest that a $M_{\bullet}\gtrsim{1000}\>\mathrm{M_{\odot}}$ GW merger with similar IMBH masses is a characteristic GW fingerprint of star cluster assembly with simultaneous SMBH seed formation at high redshifts. At lower redshifts the IMBH formation and GW mergers may still proceed in dense, low metallicity environments. Besides GWs, tidal disruption events with TDE rates up to $\Gamma\sim 4$$5\times 10^{-5}$ yr-1 per assembling star cluster might reveal a population of such low redshift IMBHs in the upcoming extragalactic wide field time domain surveys.
@@ Acknowledgments
The numerical simulations were performed using facilities in Germany hosted by the Max Planck Computing and Data Facility (MPCDF) and the JUWELS Booster of the Jülich supercomputing centre (GCS project 59949 frost-smbh-origins). TN acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence "ORIGINS". The development of the SEVN code was enabled by M. Mapelli’s ERC Consolidator grant DEMOBLACK by the European Research Council under contract no. 770017.